# A Typology of Synthetic Datasets for Dialogue Processing in Clinical Contexts

### v2

@@ 4.2.  On Beyond Microdata
There are also other factors that play important roles in the “realism” of a dataset. Recall the MTS-Dialog dataset of ben-abacha-etal-2023-empirical, which consists of “dialogues” that were imagined and “screen-written” de novo by humans (in this case, trained clinicians) based on short clinical case vignettes. Now, imagine a hypothetical second dataset (“dataset B”) in which, instead of being screen-writiten dialogue, the discourse samples were in fact the result of transcribed conversations taking place between two actors, posing as a doctor-patient dyad and improvisationally acting out a scenario based on the same case vignettes; the dataset described by fareez2022dataset follows this method of creation, as does the aci subset of the Aci-bench dataset (yim2023aci).
In both of these examples, we would consider the resulting dataset to be synthetic, in the sense that neither one reflects the language produced by actual “real world” participants in the true interaction that the dataset is purporting to represent (i.e., a clinical encounter). Our typology in its present form would categorize both as Type 3 datasets along their human dimensinos, [^9] As the microdata are being “made up” (as opposed to having naturally arisen or being modified in situ). and in both cases we can safely say that the context that gave rise to the linguistic information contained in the dataset is surely more artificial than in a hypothetical third dataset (“dataset C”) containing “genuine” conversations between patients and providers captured “in the wild” (a Type 1 dataset, according to our typology).
@@
Depending on the quality of the machine translation system, the resulting dataset may be “equivalent” in terms of its lexical and syntactic representativeness, and may also be equivalently useful in terms of its semantic content (i.e., the set of clinical conditions and events being described). [^10] We note that the extent to which this is true will depend heavily on the specifics of its intended environment of use. If our hypothetical French-speaking clinical population is in France itself, the difference may be relatively small. If it is meant for use elsewhere in the Francophone world, there may be specific medical concepts and events that are missing or over-represented. In such a situation, we would argue that the translated dataset was of limited external validity. In other words, the impact of the synthesis may be minimal along those dimensions. However, the pragmatic and discursive content may be wildly inappropriate, as norms of clinical communication and patient expectations vary significantly between cultures and countries, and French providers and patients will not speak the same way to one another as their American counterparts would. [^11] We note that, even within societies, norms of clinical communication vary greatly across numerous social and demographic factors, as well as between clinical specialties. This is one of the reasons why we are particularly wary of synthetic data in the context of clinical dialogues. As such, even assuming a very high-quality translation, the dataset is likely no longer as valid, in terms of its linguistic representativeness, as it may have been before being made synthetic. This is likely the case to a degree that is much larger than might be expected by a linguistically naïve researcher equipped with an MT algorithm. [^12] Despite technical advances, translation remains, of course, notoriously challenging (Anonymous:1968p9217) and finicky (Carlson:2007p3777). Clinical contexts bring numerous specialized challenges for the use of MT (mehandru_2022).
@@
Additionally, consider that the extra-linguistic factors represented in our hypothetical dataset may be severely impacted by this sort of perturbation. In this particular example, the translated conversation may indeed be adequately representative in terms of its linguistic and clinical content, but its contextual representativeness may be impaired. A patient-provider discussion about the management of end-stage kidney disease will include discussion of lab values, diet, etc., all of which may translate into a French context with adequate fidelity. However, the parts of the conversation involving the logistics of insurance coverage for dialysis, or of organizing a social worker to assist in transportation to and from dialysis sessions, would differ wildly between American and French contexts.
@@
This vignette is inspired by an actual example we encountered during our literature review, in which Aci-bench (a Human Synthetic Type 3 dataset) was used in a one-shot setting to prompt LLMs to generate equivalent dialogues in the Najdi Arabic dialect, with the goal of creating a dataset for use in cultural adaptation of an LLM almutairi-etal-2024-synthetic. According to our typology, the resulting data is machine-generated constrained by human-synthetic and falls under Human Type 3, Machine Type 3. As LLMs are increasingly deployed at scale, hybrid synthetic datasets of this type, will likely become more common. However, the validity and representativeness of such datasets must be carefully assessed.
@@ 5.  Conclusion
Although synthetic data sets are widely used for dialogue generation, the boundaries of what makes a dataset synthetic are not always clear and they vary among fields. Focusing on the healthcare domain and only on dialogue data, we propose a typology to clarify the types of synthetic data sets in terms of the human and machine interventions.
Academic literature on synthetic data focuses heavily on the mechanics of how synthetic data elements are generated, and our typology focuses instead on the role that the synthesis plays in the overall “shape” of the dataset. Our typology provides authors with the ability to speak in a more clear and granular way about what aspects of their dataset are syntheetic, to what degree, and from what source (human vs. machine).
Our literature review revealed a clear pattern of increased prevalence of fully-synthetic (Type 3) datasets, particularly of machine-generated dialogues following the widespread availability of LLMs in 2023. Given this, our works’s broader perspective on syntheticity will help researchers better assess and contextualize the validity and representativeness of these datasets.
The relationship between synthesis and a dataset’s larger contextual factors remains currently unexplored. As future work, we hope to expand our work to capture realism and synthesis along non-linguistic and contextual factors. [^13] The extent to which there exists a valid a dichotomy between “linguistic” and “non-linguistic” content in a dataset is surely up for debate.
@@ 6.  Limitations
Currently, the ability to characterize amounts of synthesis along these dimensions is limited, and at the present time we consider these questions to be left for future work.

### v3

@@ 4.2.  On Beyond Microdata
There are also other factors that play important roles in the “realism” of a dataset. Recall the MTS-Dialog dataset of ben-abacha-etal-2023-empirical, which consists of “dialogues” that were imagined and “screen-written” de novo by humans (in this case, trained clinicians) based on short clinical case vignettes. Now, imagine a hypothetical second dataset (“dataset B”) in which, instead of being screen-written dialogue, the discourse samples were in fact the result of transcribed conversations taking place between two actors, posing as a doctor-patient dyad and improvisationally acting out a scenario based on the same case vignettes; the dataset described by fareez2022dataset follows this method of creation, as does the ACI subset of the ACI-bench dataset (yim2023aci).
In both of these examples, we would consider the resulting dataset to be synthetic, in the sense that neither one reflects the language produced by actual “real world” participants in the true interaction that the dataset is purporting to represent (i.e., a clinical encounter). Our typology in its present form would categorize both as Type 3 datasets along their human dimensions, [^8] As the microdata are being “made up” (as opposed to having naturally arisen or being modified in situ). and in both cases we can safely say that the context that gave rise to the linguistic information contained in the dataset is surely more artificial than in a hypothetical third dataset (“dataset C”) containing “genuine” conversations between patients and providers captured “in the wild” (a Type 1 dataset, according to our typology).
@@
Depending on the quality of the machine translation system, the resulting dataset may be “equivalent” in terms of its lexical and syntactic representativeness (due to the similarity between English and French), and may also be equivalently useful in terms of its semantic content (i.e., the set of clinical conditions and events being described). In other words, the impact of the synthesis may be minimal along those dimensions. However, the pragmatic and discursive content may be inappropriate, as norms of clinical communication and patient expectations vary significantly between cultures and countries. [^9] We note that, even within societies, norms of clinical communication vary greatly across numerous social and demographic factors, as well as between clinical specialties. This is one of the reasons why we are particularly wary of synthetic data in the context of clinical dialogues. As such, even assuming a very high-quality translation, the dataset is likely no longer as valid, in terms of its linguistic representativeness, as it may have been before being made synthetic. This is likely the case to a degree that is much larger than might be expected by a linguistically naïve researcher equipped with an MT algorithm. [^10] Despite technical advances, translation remains, of course, notoriously challenging (Anonymous:1968p9217) and finicky (Carlson:2007p3777). Clinical contexts bring numerous specialized challenges for the use of MT (mehandru_2022).
@@
Additionally, consider that the extra-linguistic factors represented in our hypothetical dataset may be severely impacted by this sort of perturbation. In this particular example, the translated conversation may indeed be adequately representative in terms of its linguistic and clinical content, but its contextual representativeness may be impaired. A patient-provider discussion about the management of end-stage kidney disease in American English will include discussion of lab values, diet, etc., all of which may translate into a French context with adequate fidelity. However, the parts of the conversation involving the logistics of insurance coverage for dialysis, or of organizing a social worker to assist in transportation to and from dialysis sessions, would differ wildly between the two contexts.
@@
This scenario is inspired by an actual example we encountered during our literature review, in which ACI-bench (a Human Synthetic Type 3 dataset) was used in a one-shot setting to prompt LLMs to generate equivalent dialogues in the Najdi Arabic dialect, with the goal of creating a dataset for use in cultural adaptation of an LLM almutairi-etal-2024-synthetic. According to our typology, the resulting data is machine-generated while also being originated by a human-driven de novo synthesis process and falls under Human Type 3, Machine Type 3. As LLMs are increasingly deployed at scale, hybrid synthetic datasets of this type will likely become more common. However, the validity and representativeness of such datasets must be carefully assessed.
@@ 5.  Conclusion
Although synthetic datasets are widely used for dialogue generation, the boundaries of what makes a dataset synthetic are not always clear, and they vary among fields. Focusing on the healthcare domain and only on dialogue data, we propose a typology to clarify the types of synthetic datasets in terms of the human and machine interventions.
Academic literature on synthetic data focuses heavily on the mechanics of how synthetic data elements are generated, and our typology focuses instead on the role that the synthesis plays in the overall “shape” of the dataset. Our typology provides authors with the ability to speak in a more clear and granular way about what aspects of their dataset are synthetic, to what degree, and from what source (human vs. machine).
Our literature review revealed a clear pattern of increased prevalence of fully-synthetic (Type 3) datasets, particularly of machine-generated dialogues following the widespread availability of LLMs in 2023. Given this, our work’s broader perspective on syntheticity will help researchers better assess and contextualize the validity and representativeness of these datasets.
The relationship between synthesis and a dataset’s larger contextual factors remains currently unexplored. As future work, we hope to expand our work to capture realism and synthesis along non-linguistic and contextual factors. [^11] The extent to which there exists a valid dichotomy between “linguistic” and “non-linguistic” content in a dataset is surely up for debate.
@@ 6.  Limitations
Currently, the ability to characterize amounts of synthesis along these dimensions is limited, and at the present time, we consider these questions to be left for future work.
Acknowledgments
Sergiu Nisioi is supported by the project “Romanian Hub for Artificial Intelligence - HRIA”, Smart Growth, Digitization and Financial Instruments Program, 2021-2027, MySMIS no. 351416. Steven Bedrick is supported by the National Library of Medicine of the National Institutes of Health under award number 5R01LM011934-10 (“Semi-structured Information Retrieval in Clinical Text for Cohort Identification”).
@@ 7.  Bibliographical References
Appendix A What is a Dataset?
In response to the complexities surrounding the question of whether a given dataset is “synthetic” or not, we propose a conceptual framework whose aim is to provide structure and clarity when describing datasets that have a synthetic component. To facilitate meaningful comparisons, we must first define what we mean by “dataset,” and propose here a broad and “teleological” definition.We set as a foundation that a dataset is a finite collection of information-bearing objects, in the sense described by bucklandInformationThing1991. Such a collection may be either digital or physical in nature. Note that we refer here to a “collection” as opposed to a “grouping”, “assembly”, “set”, etc. Not all groupings of information objects constitute datasets. A random and incidental set of bits would not constitute a dataset, nor would a random pile of books encountered on the roadside.
A collection implies a collector (i.e., an external agent with at least some degree of intentionality behind its assembly). Furthermore, a collection necessarily brings with it some notion of an intended user whose needs or goals the collection is meant to support. [^12] The user and the collector may, of course, be the same entity. Another key property of a collection is that it involves some kind of extrinsic criteria for what is and is not included in the collection. [^13] In the context of traditional library and archival collections, questions of ownership and accessibility also play into the definition of a collection (Lee:2000).
This definition, however, is insufficient. An intentionally collected set of bits representing random numbers may not on its own constitute a dataset. Therefore, we introduce a second constraint to our definition. Our collection of information-bearing objects must be documents in the sense meant by briet1951, who defined a document as “any concrete or symbolic indication, preserved or recorded, for reconstructing or for proving a phenomenon, whether physical or mental.”141414“Tout indice concret ou symbolique, conservé ou enregistré, aux fins de représenter ou de prouver un phénomène ou physique ou intellectual.”, briet1951 cited in bucklandInformationThing1991.
According to this definition, a crucial distinction between “a collection of random bits” and a “dataset” is that the latter has an underlying intention of some sort on the part of whoever is identifying it as such (its creator or collector), and that this intention is reflected in the nature of the dataset’s contents. The “concrete or symbolic indication” is being intentionally “preserved or recorded” for some purpose. Thus, we can conclude that it is necessary for that purpose that it be a meaningful and representative representation. Otherwise it would be useless for “reconstructing or … proving a phenomenon.” [^15] In the case of a dataset comprising random noise, the intended point of the dataset is that the noise is in fact random according to some particular definition or scheme, and the existence of the dataset would constitute an assertion of that property on the part the dataset’s authors.
Under this part of our definition, we argue that part of what makes a dataset a dataset is that it is intended by its originators to be representative of “something”, hence, our definition’s description as “teleological.” What that “something” is may vary wildly, as may the degree to which a given dataset actually represents whatever it is meant to represent. These are important questions, but even a dataset that is flawed, incomplete, broken, etc., still counts as a dataset.
To illustrate this distinction, consider the bits representing the set of digital music files on one of the authors’ laptops. This is comprised of a large grouping of binary files and associated metadata, but taken on its own, stripped of context and intention, it is simply that: a grouping. It serves no purpose and has no intrinsic meaning, other than to exist and be stored in his computer’s filesystem. Its constituent bits may be interpretable by this or that piece of software, but are essentially the digital equivalent of a random and incidental pile of books.
In this particular case, it is reasonable to argue that they represent a collection, in the sense previously described. They have been intentionally collected, in that they have been purposefully assembled over a period of time as new music has been added, and there also exist numerous possible digital music files that are not in the collection. However, it would be nonsensical to say that this particular collection is or is not “representative”, “meaningful”, “useful”, “complete”, “well-formed”, etc.: each of those descriptors implies a telos of some kind. In other words, each immediately prompts the questions, “for what?” and “for whom?” “Meaningful” — to whom, in what way, and in what context? The files in their present form lack any such thing — they are simply files.
Now, imagine that the author in question decides to try and use this collection as part of an analysis, perhaps to visualize his music consumption patterns. We argue that it is at this point, and not before, that our collection of information objects becomes a dataset: the point at which it is first intentionally collected in order to represent, reconstruct, prove, etc. a particular phenomenon. In other words, the point at which the bits become part of a process of problem formulation. [^16] See Passi:2019 for an excellent discussion on this topic, including on the contingent and bidirectional relationship between data and problem formulation.
Next, we add an additional constraint: in addition to the intention on the part of somebody to use a collection for some particular purpose, we also argue that there must also be an intention for some other party to understand the collection as being similarly meaningful. This is closely linked to the notion of “audience” in our discussion of collections above, but is more specifically linked to the telos of the dataset in question.
At this point, our definition has a “what” (collection of information objects), a “when” (the point at which somebody treats the collection as being a meaningful representation of a particular phenomenon), and a “who” (whatever additional audience or user community the author expects to align with them as to the nature and purpose of the dataset). In the interests of theoretical completeness, we add one more constraint, which may be categorized under the heading of “whether,” [^17] In the descriptive sense of “whether or not it is possible”, not the prescriptive sense of “whether or not it is a good idea”. and argue that for a collection to be considered a dataset, it must also contain information objects that in terms of their structure and content could plausibly be used from a functional or operational standpoint to achieve the dataset’s purpose.
In other words, if the point of the dataset is to visualize the author’s music consumption patterns, it must necessarily be comprised of documents containing the appropriate kinds of information (e.g., digital files containing music, metadata recording listening patterns, financial records of purchases, etc.), as opposed to being (e.g.) made up of random bits, pictures of cats, etc. This part of our definition may at first appear somewhat tautological. However, we include it to help rule out counter-examples and clarify the functional nature of our definition. [^18] We express our thanks to Dr. Colin Koopman for this part of our definition (koopman_2025). An analogy with measurement theory (brewer_crano_2014) is appropriate. This part of our definition attempts to rhyme more with “face validity” than it does with “construct validity” or “measurement validity”, in that we are not referring to how well the datasets’ contents serve its intended purpose but rather are speaking about basic functional plausibility.
Nothing in our definition means that a dataset built to serve one telos may not be used in pursuit of another, or that a dataset may not be designed in pursuit of multiple telea. It does, however, help make apparent the fact that re-purposing of data is a less straightforward undertaking than it often appears and requires careful consideration. Additionally, our definition has nothing to say about what kinds of purposes are or are not “valid,” or limits to how general or specific the purpose of a dataset may be.
Appendix B Search Strategies
B.1.  PubMed Search Strategy
Our PubMed search strategy consisted of three separate searches, with the goal of identifying published datasets of clinical dialogues, either synthetic or natural. This was conducted as a supplement to our preliminary literature search conducted using less precise methods, and should be considered current as of April, 2025.
The first search used solely keywords, and attempted to directly capture papers reporting the creation of datasets relating to dialogues, and was as follows: (((patient AND provider) or (clinical)) AND (conversation* OR dialog OR dialogue)) AND (corpus OR dataset)
The second attempted to leverage MeSH index terms, and did not specifically include terms relating to datasets or corpora, but did add a handful of keywords to reduce scope: ("Artificial Intelligence"[MeSH Terms] OR "Machine Learning"[MeSH Terms] OR "Natural Language Processing"[MeSH Terms] OR "Information Storage and Retrieval"[MeSH Terms] OR "Pattern Recognition, Automated"[MeSH Terms]) AND ("Physician-Patient Relations"[MeSH Terms] OR "Referral and Consultation"[MeSH Terms] OR "Interviews as Topic"[MeSH Terms] OR "Patient Simulation"[MeSH Terms] OR "Communication"[MeSH Terms]) AND ("conversation" OR "dialogue" OR "dialog" OR "discourse")
The third returned to a keyword approach, and focused on patient-provider conversations: (((patient AND provider) or (clinical)) AND (conversation* OR dialog OR dialogue)) AND (corpus OR dataset)
The three searches together yielded 679 unique results, of which the vast majority were false positives. An LLM (Gemma 3 21b) was used to perform further filtration. First, we instructed the model to classify each retrieved paper according to whether its abstract described “automated processing of either spoken or written clinical language in the context of patient-provider communication.” This reduced the set to 371 articles; manual inspection revealed satisfactory performance, and we were able to verify that specific known true-positive results were correctly classified by this method, further supporting its efficacy.
We, then, performed a second step of LLM-based filtration, instructing the model to classify each of those articles as to whether the abstract was “describing a paper that created a new corpus or dataset with the intention of it being re-usable by others.” The model was instructed to prioritize recall over precision.
This reduced our set of articles to review to 125. Both prompts are rendered in the boxes from Appendix D. We used similar heuristic methods to verify that this filtering method performed at a satisfactory level of accuracy. The precision was reasonably high, in that articles in this set did in general contain notable discussion of the details of the datasets that they used, and generally did all describe the creation of novel corpora. However, many were describing corpora that were outside of our review’s scope (i.e., datasets of QA pairs, etc.), or described corpora that were not releasable due to reasons of privacy or other such constraints. Only 4 papers matched our criteria and were included in our review.
B.2.  ACL Anthology Search Strategy
We implemented a script to retrieve, filter, and annotate papers from the ACL Anthology https://aclanthology.org/. The process involves scraping the HTML content of the full ACL proceedings between 2019-April, 2025 (EMNLP, ACL, EACL, NAACL, LREC) including all the workshops associated. Each paper title, abstract, authors and paper URL was extracted.
Papers were filtered based on the presence of the keyword medic in either the title or the abstract. This filtering is high recall, case-insensitive, and matches partial substrings, resulting a total of 591 papers. For each filtered paper, the corresponding full text document is extracted and submitted to the ChatPDF API. [^19] https://www.chatpdf.com/ A fixed prompt (see Appendix D) was used to request a structured JSON response with answers to the following binary and list-format questions: (1) whether a synthetic dataset was created, (2) whether the dataset contains dialogues, (3) whether human evaluation was conducted, (4) whether the dataset was used in a downstream task. We further reduce the list to 180 papers based on criteria (1) and (2). A manual review of the titles and abstracts further narrowed the set to 23 papers. Full-text examination of these 23 papers resulted in 11 that matched all criteria of interest and were included in our final review.
B.3.  dblp Search Strategy
The search employed on https://dblp.org was broad and covered a range of related fields, including speech processing, medical informatics, dialogue systems, artificial intelligence, argumentation, and information retrieval. To ensure high recall, 202020(med* | syn* | sim*) & dial* the search targeted titles containing words beginning with med* (e.g., medical, medication), syn* (e.g., synthesis, synthetic), or sim* (e.g., simulation, simulating) in combination with dialo* (e.g., dialogue).
To refine the results, a filter was applied to exclude publications from major venues associated with the Association for Computational Linguistics (ACL) and preprints. The final corpus was restricted to articles published in selected conferences, workshops, and journal series, [^21] In total 362 streamid’s corresponding to different publications. including INTERSPEECH, EUROSPEECH, ICASSP, IJCAI, SIGIR, AAAI, MEDINFO, the Journal of Biomedical Informatics, BMC Medical Informatics and Decision Making, Computers in Biology and Medicine, CEUR-WS, FAIA, CCIS, SCI, and relevant Springer proceedings, among others.
In total, 356 papers were identified through this search. Based on title screening, 45 of the retrieved papers were selected for further manual review of abstract and content. Among these, only 10 papers have been marked as suitable for our review.
Appendix C Synthetic Datasets
Source Initial Hits After First Filter After Review Included
PubMed 679 371 (LLM: clinical language) 125 (LLM: dataset creation) 4
ACL Anthology 591 180 (LLM: dataset+dialogue) 23 11
dblp 356 - 45 10
Table 2: Search strategy summary. The first filter for PubMed and ACL Anthology are based on LLM prompting. There is an overlap between the three search strategies and the total number of unique papers is 20.
Dataset Acronym Language Types
ben-abacha-etal-2023-empirical MTS-DIALOG en
Human Type 3, Machine Type 1
liu-etal-2021-towards ESConv en
Human Type 3, Machine Type 1
fareez2022dataset en
Human Type 3, Machine Type 2
papadopoulos-korfiatis-etal-2022-primock57 PriMock57 en
Human Type 3, Machine Type 2
yim2023aci ACI-bench en
Human Type 3, Machine Type 2
almutairi-etal-2024-synthetic ar-sa
Human Type 3, Machine Type 3
zheng-etal-2023-augesc AugESC en
Human Type 3, Machine Type 3
Chen:2023c IMCS-21 zh
Human Type 2, Machine Type 1
shi-etal-2023-midmed MidMed zh
Human Type 2, Machine Type 2
zeng-etal-2020-meddialog MedDialog zh / en
Human Type 1, Machine Type 1
liu2022meddg MedDG zh
Human Type 1, Machine Type 2
li2021semi KaMed zh
Human Type 1,Machine Type 2
xu-etal-2024-reasoning EMULATION zh
Human Type 1, Machine Type 2
binici2024medsage MEDSAGE ch
Human Type 1, Machine Type 2
zhang-etal-2023-huatuogpt HuatuoGPT zh
Human Type 1, Machine Type 3
nishiyama-etal-2024-assessing ja
Human Type 1, Machine Type 3
han-etal-2024-psydial-personality PSYDIAL ko
Human Type 1, Machine Type 3
chen2023investigating DoPaCo en
Human Type 1, Machine Type 3
wang-etal-2024-notechat NoteChat en
Human Type 1, Machine Type 3
moser2024synthetic en, de
Human Type 1, Machine Type 3
Table 3: The list of identified synthetic datasets and their categories.
Appendix D LLM Prompts
Gemma 3 27b: clinical language filter Your role is to review an abstract of a journal article, and decide whether it should be included in a literature review. We are looking for articles about automated processing of either spoken or written clinical language in the context of patient-provider communication. Examples would include: clinical conversations, discourse, written narrative communications, etc. Either transcribed or recorded language is acceptable. If you are not sure, or if the article is borderline, please consider it in-bounds and err on the side of inclusion.
Gemma 3 27b: dataset creation filter Your role is to review an abstract of a journal article, and determine whether or not the abstract is describing a paper that created a new corpus or dataset with the intention of it being re-usable by others. Possible things to look for: - The abstract might explicitly mention that the dataset produced as part of the article was released or otherwise made publicly available. - The abstract might talk about how the authors developed a novel corpus, specifically because of a lack of already-existing data or because existing data was insufficient in some way. Remember, all of these papers will have involved making a dataset or corpus of some kind. We are looking specifically for papers that where one of the final products of the research, or where the main point of the research, was to create a new dataset or corpus, especially if it is described as being released or made available. Simply mentioning an already-existing dataset is not enough.
ChatPDF: dataset+dialogue Provide a json response using the following structure: { "synthetic_dataset_created": true/false, "contains_dialogues": true/false, "human_evaluation": true/false, "uses_dataset_in_downstream_task": true/false } To answer the following questions: - Do the authors create a synthetic dataset in this paper? - Does it contain dialogues? - Does it have a human evaluation of the dataset? - Does it use the dataset in a downstream task?